Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET Sample Environment SPB_mig
Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 390     Entry time: 30 Jan 2023, 09:49
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 30.01.2023 
General Announcement:
-	 

CTRL:
- Nothing to report.

DA:
- DA-OCD service has started from this week.

EEE:
- PLC had an OCD call on Saturday (motor SA3 M2). XRO could continue and further investigation is ongoing.
- SEPS to be done today in SA3.

VAC:
- Nothing to report.

XPD:
- Minor HIREX issue on the weekend but did not affect activities.

DET:
- Jungfrau detector had a small bug in Karabo control device. The device has to be re-instantiated when detector is rebooted. CTRL is aware and preparing fix.

ITDM: 
- Nothing to report.

XRO:
- On SA3 M2: issue related to PLC update in December; apparently resulted with loss of calibration in one motor. All alignment were done with other parameters based on beam reflection. 
- Steffen Hauf comments there was two different issues with different motors: 1. the calibration was solved end of last week; 2. from Saturday it was a problem with encoder which will be fixed this week 
with EEE. 

LAS:
- Stages with problem but running.
- SA1 will be ready for users by the end of today.

UNSYS:
- Small issue with SA1/cell22 air coils were fixed by Suren K.

SPB:
- Thanks to supporting groups for the bring-up days.

FXE:
- Nothing to report.

MID:
- Question on the availability of PP-laser in SA2: Daniel K. informs that is expected by the end of today but this will be confirmed with other LAS colleagues.
- Question about intensity in SA2: A. Galler informs there is no news on the cause of low intensity; Raul V. as PRC mentions currently it was reached about 500uJ at 9keV.

HED:
- Nothing to report.

SCS:
- Bring-up days today and SEPS interlock tests.

SQS:
- Nothing to report.
ELOG V3.1.4-7c3fd00